OpenStudy (anonymous):

-4a+9 greater than or equal to 11a

myininaya (myininaya):

-4a+9>=11a -4a-11a>=-9 -15a>=-9 a<=-9/-15=3/5

myininaya (myininaya):

a<=3/5 (-inf,3/5]

OpenStudy (anonymous):

can someone view my question

OpenStudy (anonymous):

x is less than 3/5 and jamba juice is delicious.

OpenStudy (anonymous):

Thanks everyone who answered my question :)